student A prepared 0.044 M concentration of the stock sodium salicylate solution and using a Mohr pipet transferred 5ml of the 0.044 M sodium salicylate stock solution into a 100 ml volumetric flask labelled as sample 1. the solution was diluted to the 100 ml mark with the acidified 0.02 M iron 3 chloride solution. what is the concentration of iron 3 salicylate complex in the sample 1 flask.
1
Expert's answer
2020-10-08T13:54:58-0400

Equation of reaction

3C7H5NaO3 +FeCl3----------->C21H15FeO9+3NaCl


C7H5NaO3 : C21H15FeO9 mole ratio 3:1

Moles of C7H5NaO3 =(0.044÷1000)×5=0.00022

Using the mole ratio, the moles for C21H15FeO9 will be (0.00022÷3)×1=0.000073moles

This is present in 100ml. To find concentration in 1L,

(0.000073÷100)×1000=0.00073

=7.3×10-4mols L-1